OpenStudy (anonymous):

the volume of a rectangular prism is 2,058 cubic cm. the length of the prism is 3 times the width. the hieght is twice the width. find the width of the prism.

jhonyy9 (jhonyy9):



jhonyy9 (jhonyy9):

l=3w h=2w area of base =l*w volum of prism = area of base * height 2,058 = l*w *h but you know that l=3w and h=2w so than we get 2,058 = 3w*w*2w so 2,058 = 6w^3 divide both sides by 6 0,343 = w^3 w = cuberoot(0,343)